OpenStudy (anonymous):

1/5-x=3 2.8 -2.8 -3.2 3.2

OpenStudy (anonymous):

everyone wait

OpenStudy (anonymous):

niti, do u know how to solve a linear equation>

OpenStudy (anonymous):

man, i am in the middle of the test i don't have time to solve i have another 20 questions :( !!!!

OpenStudy (anonymous):

isn't that the point of the test?

OpenStudy (anonymous):

sorry, if you want to help me, okay if not, than thank you :) !

OpenStudy (anonymous):

-2.8

OpenStudy (anonymous):

thanks, you saved my life !!!!

OpenStudy (anonymous):

do all ppl there do this paul>??

OpenStudy (anonymous):

can you tell me what you did pleaes him?

OpenStudy (anonymous):

i helped someone cheat?

OpenStudy (anonymous):

ha, no not that, i have to understand linear equations... and i dont really do that currently

OpenStudy (anonymous):

well sometimes we have to..lol

OpenStudy (anonymous):

0.2 - x =3 -x = 3 - 0.2 = 2.8 x=-2.8

OpenStudy (anonymous):

i know these things, ppl but i have around 80 questions, and i have only 60 minutes to do this test, and all this hall is with students and teachers around !!

OpenStudy (anonymous):

oh good.

OpenStudy (anonymous):

so you just cheat right in front of everyone

OpenStudy (anonymous):

thats guts for you

OpenStudy (anonymous):

we can use our laptops, if we don't know something..

OpenStudy (anonymous):

hahaha....open book test..thts wht we call it..ull never learn lyk this

OpenStudy (anonymous):

f o

OpenStudy (anonymous):

1/5-x=3 -x= 3-(1/5) - x=(15-1)/5 - x = 14/5 x= -14/5 x =-2.8